Fermi National Laboratory


Accelerator Update for November 21, 2005 - November 23, 2005

Monday November 21

The day shift began with the Tevatron (TeV) in shot setup, with Operations monitoring the Antiproton Source (Pbar) stack, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, with LCW experts working on the water system for Pbar power supplies, and with MiniBooNE and NuMI taking beam.

Operations established store 4515 at 8:10 AM with an initial luminosity of 136.8E30.

Operations established beam to the Meson MTest and MCenter experiments at 8:22 AM.

Pbar resumed stacking at 8:30 AM.

Operations prepared the Recycler to stash antiprotons at 10:09 AM. The Recycler took 6 transfers. Pbar resumed stacking 12:46 PM. MTest resumed taking beam.

MCenter resumed taking beam at 1 PM.

Main Injector (MI) experts began a RF study at 2:32 PM.

At 4:16 PM, the TeV suffered a multi-house quench due to a sector B11 separator spark. Store 4515 was lost.

At 4:39 PM, the NuMI beam permit tripped off, probably due to MI RF problems. MIRF13 was off for the study and then MIRF5 developed a LCW leak and had to be turned off for repairs.

Operations began preparing the TeV for an access at 5:15 PM.

The MI RF study ended at 5:34 PM. They have to reconfigure the RF stations for normal running. TeV expert reported at 5:53 PM that they found super insulation sticking out of the blow-off plates on the sector B17 spool piece. Cryo system technicians will begin warming up the B1 cryo system.

B17 Spool Piece Surrounded by Insulation

A TeV Expert Holds Up a Streamer

Cryo techs reported that they were also going to warm up the TeV sector A4 area to inspect a leaking lead at 6:24 PM.

A Recycler stash, which had probably started sometime around 7 PM, ended at 8:17 PM. Pbar, MTest, and MCenter resumed normal operations by 8:30 PM.

At 11:43 PM, the P150 transfer line suffered a vacuum burst. The vacuum pumps reset right away, so there was little downtime from this.

Tuesday November 22

The midnight shift began with TeV experts waiting for the sector B1 cryo warm up, with Operations monitoring the stash, with Pbar stacking, and with MiniBooNE, NuMI, MTest, and MCenter taking beam.

Operations prepared CDF for a collision hall access at 8:27 AM.

A MI expert began a coalescing study at 1:40 PM.

A Pbar expert reported at 3:47 PM a problem with DRF2.

At 6:12 PM, Operations reported that QXR had stopped working. An operator contacted an expert. This problem halted beam to Switchyard (SY).

A controls expert restored beam to SY at 8:15 PM by initializing the MI beamline proton transfer parameter.

Wednesday November 23

The midnight shift began with cryo experts warming up sector B1 in the TeV, with Operations monitoring the Pbar stack, with Pbar experts conducting a reverse proton study, with Pelletron experts conducting Electron Cooling studies, and with MiniBooNE, NuMI, MTest, and MCenter taking beam.

At 12:38 AM, the P1 transfer line suffered a vacuum burst. The vacuum recovered and beam returned to Pbar and SY.

The Electron Cooling studies ended at 1:32 AM.

At 5:25 AM, Operations reported a problem with the Booster tune. Operators investigated but couldn’t get better than a 50% efficient beam. At 6 AM, Operations called in an expert.

Reverse proton studies ended at 7:02 AM.

A Booster expert brought the Booster back to normal efficiency at 7:45 AM.

The Plans for Wednesday

The plans for today are to continue with the warm up of TeV sector B1, which should occur on Thursday. Experts will be able to investigate the damage to the bypass on Friday morning. We already know that the spool piece needs to be replaced, which means the TeV would be down for about a week. If the bypass must be replaced, that will add to the TeV downtime.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 11/23/2005   email Fermilab